OpenStudy (anonymous):

Use the double-angle formulas to rewrite the following: 5-10sin^2x

OpenStudy (tkhunny):

Well, do it. \(\cos(2x) = \cos^{2}(x) - \sin^{2}(x) = 2\cos^{2}(x) - 1 = 1 - 2\sin^{2}(x)\) Multiply that whole thing by 5 and see what comes out.

OpenStudy (anonymous):

\[\cos 2x=\cos ^2x-\sin ^2x=1-\sin ^2x-\sin ^2x=1-2\sin ^2x\]